OpenStudy (gamenerd123):

Solve for y. −140=18+4(5y−2)

RhondaSommer (rhondasommer):

\[−140=18+4(5y−2)\\-140=18+20y-8\\150=20y\\y=7.5\]

OpenStudy (gamenerd123):

Thx bro